### v1

@@ Abstract
keywords:
Structural Pruning, Visual Prompt, Sparse Neural Network1 Introduction
Recently, large-scale neural networks, particularly in the field of vision and language modeling, have received upsurging interest due to the promising performance for both natural language [1, 2, 3] and vision tasks [4, 5]. While these models have delivered remarkable performance, their colossal model size, coupled with their vast memory and computational requirements, pose significant obstacles to model deployment. To solve this daunting challenge, model compression techniques have re-gained numerous attention [6, 7, 8, 9, 10, 11] and knowledge distillation can be further adopted on top of them to recover optimal performance [12, 13, 14]. Among them, model pruning is a well-established method known for its capacity to reduce model size without compromising performance [15, 16, 17] and structural model pruning has garnered significant interest due to its ability to systematically eliminate superfluous structural components, such as entire neurons, channels, or filters, rather than individual weights, making it more hardware-friendly [18, 19, 20, 21].
In the context of structural pruning for vision models, the paramount task is the estimation of the importance of each structure component, such as channel or filters. It is a fundamental challenge since it requires dissecting the neural network behavior and a precise evaluation of the relevance of individual structural sub-modules. Previous methodologies [22, 20, 23, 24, 25] have either employed heuristics or developed learning pipelines to derive scores, achieving notable performance. Recently, the prevailingness of natural language prompts [26, 27] has facilitated an emerging wisdom that the success of AI is deeply rooted in the quality and specificity of data that is originally created by human [28, 29]. Techniques such as in-context learning [30, 31, 32] and prompting [33, 34, 35, 36, 37] have been developed to create meticulously designed prompts or input templates to escalate the output quality of LLMs. These strategies bolster the capabilities of LLMs and consistently achieve notable success across diverse downstream tasks. This offers a brand new angle for addressing the intricacies of structural pruning on importance estimation of vision models: How can we leverage the potentials within the input space to facilitate the dissection of the relevance of each individual structural component across layers, thereby enhancing structural sparsity?
One straightforward approach is directly editing input through visual prompt [38] to enhance the performance of compressed vision models [39]. The performance upper bound of this approach largely hinges on the quality of the sparse model achieved by pruning, given that prompt learning is applied post-pruning. Moreover, when pruning is employed to address the intricate relevance between structural components across layers, the potential advantages of using visual prompts are not taken into consideration.
@@
Moreover, the complexities associated with inherent channel dependencies render the generation of sparse channel masks a challenging task. Due to this reason, many previous arts of pruning design delicate pruning metrics to recognize sparse subnetworks with smooth gradient flow [40, 41, 42]. To better handle the channel dependencies across layers during channel pruning, we propose to learn sparse masks using a recurrent mechanism. Specifically, the learned sparse mask for the recent layer largely depends on the mask from the previous layer in an efficient recurrent manner, and all the masks are learned by incorporating the extra information provided by visual prompts. The PASS framework is shown in Figure 1. Our contributions are summarized as follows:
@@ 2 Related Work
Structural Network Pruning. Structural pruning achieves network compression through entirely eliminating certain superfluous components from the dense network. In general, structural pruning follows three steps: (i) pre-training a large, dense model; (ii) pruning the unimportant channels based on criteria, and (iii) finetuning the pruned model to recover optimal performance. The primary contribution of various pruning approaches is located in the second step: proposing proper pruning metrics to identify the importance of channels. Some commonly-used pruning metric includes but not limited to weight norm [43, 44, 45], Taylor expansion [17, 46], feature-maps reconstruction error [47, 48, 49, 50], feature-maps rank [51], KL-divergence [52], greedy forward selection with largest loss reduction [53], feature-maps discriminant information [54, 55, 56].
@@
Prompting. In the realm of natural language processing, prompting has been acknowledged as an effective strategy to adapt pre-trained models to specific tasks [57]. The power of this technique was highlighted by GPT-3’s successful generalization in transfer learning tasks using carefully curated text prompts [1]. Researchers have focused on refining text prompting methods [58], [59] and developed a technique known as Prompt Tuning. This approach involves using prompts as task-specific continuous vectors optimized during fine-tuning [60], [61], [36], offering comparable performance to full fine-tuning with a significant reduction in parameter storage and optimization. Prompt tuning’s application in the visual domain has seen significant advancement recently. Pioneered by Bahng et al. [62], who introduced prompt parameters to input images, the concept was expanded by Chen et al. [35] to envelop input images with prompt parameters. [38] took this further, proposing visual prompt tuning for Vision Transformer models. Subsequently, Liu et al. [63], Zheng et al. [64], Zhang et al. [65] designed a prompt adapter to enhance these prompts. Concurrently, Zang et al. [66], Zhou et al. [67, 68] integrated visual and text prompts in vision-language models, boosting downstream performance.
@@
Hypernetwork. Hypernetworks represent a specialized form of network architecture, specifically designed to generate the weights of another Deep Neural Network (DNN). This design provides a meta-learning approach that enables dynamic weight generation and adaptability, which is crucial in scenarios where flexibility and learning efficiency are paramount. Initial iterations of hypernetworks, as proposed by Zhang et al. [69], Galanti and Wolf [70], David et al. [71], Li et al. [72], were configured to generate the weights for an entire target DNN. While this approach is favorable for smaller and less complex networks, it constrains the efficacy of hypernetworks when applied to larger and more intricate ones. To address this limitation, subsequent advancements in hypernetworks have been introduced, such as the component-wise generation of weights [73, 74, 75] and chunk-wise generation of weights [76]. Diverging from the initial goal of hypernetworks, our work employs them to fuse visual prompts and model information for generating sparse channel masks.
@@ 3 PASS: Visual Prompt Locates Good Structure Sparsity
Notations. Let us consider a CNN with layers, and each layer contains its corresponding weight tensor , where {, , and } are the number of output/input channels and convolutional kernal size, respectively. The entire parameter space for the network is defined as . Similarly, a layer-wise binary mask is represented by , where “”/“” indicates removing/maintaining the associated channel. denotes our visual prompts. denotes the data of a target task.
Rationale. In the realm of structural pruning for deep neural networks, one of the key challenges is how to derive channel-wise importance scores for each layer. Conventional mechanisms estimate the channel significance either in a global or layer-wise manner [48, 43, 20, 77], neglecting the sequential dependency between adjacency layers. Meanwhile, the majority of prevalent pruning methods are designed in a model-centric fashion [20, 43, 78, 79, 22, 23]. In contrast, an ideal solution to infer the high-quality sparse mask for one neural network layer should satisfy several conditions as follows:
@@ 3.1 Innovative Data-Model Co-designs through A Recurrent Hypernetwork
Modeling the Layer Sequential Dependency. The recurrent hyper-network in PASS adopts a Long Short-Term Memory (LSTM) backbone since it is particularly suitable for capturing sequential dependency. It enables an “auto-regressive” way to infer the structural sparse mask. To be specific, the LSTM mainly utilizes the previous layer’s mask , the current layer’s weights , and a visual prompt as follows:
where the visual prompt provides an initial hidden state for the LSTM hyper-network, is the parameters of the LSTM model, and is the extra encoder to map the visual prompt into an embedding space. The channel-wise sparse masks () generated from the hyper-network are utilized to prune the weights of each layer as expressed by . represents the pruning of in-channels while denotes the pruning of out-channels.
@@
For achieving an optimal non-uniform layer-wise sparsity ratio, we adopt global pruning [81] that eliminates the channels associated with the lowest score values from all layers during each optimization step. This approach is grounded in the findings of Huang et al. [81], Liu et al. [82], Fang et al. [20], which demonstrate that layer-wise sparsity derived using this method surpasses other extensively researched sparsity ratios.
@@ 4.1 Implementation Setups
Architectures and Datasets. We evaluate PASS using four pre-trained models: ResNet-, ResNet-, ResNet- [83], and VGG- without BatchNormD [84], all pre-trained on ImageNet-1K [85]. Our evaluation contains six target tasks: Tiny-ImageNet [85], CIFAR- [86], DTD [87], StanfordCars [88], and Food [89]. The size of the inputs is scaled to during our experiments.
Baselines. We select five popular structural pruning methods as our baselines: () Group-L structural pruning [18, 20] reduces the network channels via regularization. () GrowReg [23] prunes the network channels via regularization with a growing penalty scheme. () Slim [22] imposes channel sparsity by applying regularization to the scaling factors in batch normalization layers. () DepGraph [20] models the inter-layer dependency and group-coupled parameters for pruning and () ABC Pruner [90] performs channel pruning through automatic structure search.
Training and Evaluation. We utilize off-the-shelf models from Torchvision 111https://pytorch.org/vision/stable/index.html as the pre-trained models. During the pruning phase, we employ the SGD optimizer for the visual prompt, while the AdamW optimizer is used for the visual prompt encoder and the LSTM model for generating channel masks. Regarding the baselines, namely Group-L structural pruning, GrowReg, Slim, and DepGraph, they are trained based on this implementation 222https://github.com/VainF/Torch-Pruning and ABC Prunner is trained based on their official public code 333https://github.com/lmbxmu/ABCPruner. During the fine-tuning phase, all pruned models, inclusive of those from PASS and the aforementioned baselines, are fine-tuned with the same hyper-parameters. We summarize the implementation details and hyper-parameters for PASS in Appendix C. For all experiments, we report the accuracy of the downstream task during testing and the floating point operations (FLOPs) for measuring the efficiency.
@@ 4.2 PASS Finds Good Structural Sparsity
From Figure 2, several salient observations can be drawn: ❶ PASS consistently demonstrates superior accuracy across varying FLOPs values for all six evaluated downstream tasks. On one hand, PASS achieves higher accuracy under the same FLOPs. For example, it achieves higher accuracy than baselines under FLOPs among all the datasets. On the other hand, PASS attains higher speedup444Following Fang et al. [20], we report the theoretical speedup ratios and it is defined as in achieving comparable accuracy levels. For instance, to reach accuracy levels of , , and on CIFAR, StanfordCars, and Food respectively, the PASS method consistently realizes a speedup of at least ( VS ), outperforming the most competitive baseline. This consistent performance highlights the robustness and versatility of the PASS method across diverse scenarios. ❷ In terms of resilience to pruning, PASS exhibits a more gradual reduction in accuracy as FLOPs decrease. This trend is notably more favorable when compared with the sharper declines observed in other baseline methods. ❸ Remarkably, at the higher FLOPs levels, PASS not only attains peak accuracies but also surpasses the performance metrics of the fully fine-tuned dense models. For instance, PASS excels the fully fine-tuned dense models with {, , } on CIFRAR, DTD and FOOD datasets.
@@
Superior Performance across Model Architectures. We further evaluate the performance of PASS across multiple model architectures, namely VGG- without batch normalization 555The baseline Slim [22] is not applicable to this architecture., ResNet-, and ResNet- and compare it with the baselines. The results are shown in 3. We observe that our PASS achieves a competitive performance across all architectures, often achieving accuracy close to or even surpassing the dense models while being more computationally efficient. For instance, To achieve an accuracy of on Tiny-ImageNet using ResNet-/ResNet- and accuracy using VGG-, our PASS requires fewer FLOPs compared to the most efficient baseline. These observations suggest that PASS can effectively generalize across different architectures, maintaining a balance between computational efficiency and model performance.
@@
4.3 Experiments on ImageNet and Advanced Architectures
To draw a solid conclusion, we further conduct extensive experiments on a large dataset ImageNet using advanced pre-trained models such as ResNeXt-50, Swin-T, and ViT-B/16. The results are shown in Table LABEL:Tab:AdvancedModel. We observe that our method PASS demonstrates a significant speed-up with minimal accuracy loss, as indicated by the Acc., which is superior to existing methods like SSS [91], GFP [92], and DepGraph [20]. the resulting empirical evidence robustly affirms the effectiveness of PASS across both advanced neural network architectures and large-scale datasets.
| Arch. | Method | Base | Pruned | Acc. | FLOPs |
| ResNeXt-50 | ResNeXt-50 | 77.62 | - | - | 4.27 |
| SSS [91] | 77.57 | 74.98 | -2.59 | 2.43 | |
| GFP [92] | 77.97 | 77.53 | -0.44 | 2.11 | |
| DepGraph [20] | 77.62 | 76.48 | -1.14 | 2.09 | |
| Ours (PASS) | 77.62 | 77.21 | -0.41 | 2.01 | |
| ViT-B/16 | ViT-B/16 | 81.07 | - | - | 17.6 |
| DepGraph [20] | 81.07 | 79.17 | -1.90 | 10.4 | |
| Ours(PASS) | 81.07 | 79.77 | -1.30 | 10.7 | |
| Swin-T | Swin-T | 81.4 | - | - | 4.49 |
| X-Pruner [93] | 81.4 | 80.7 | -0.7 | 3.2 | |
| STEP [94] | 81.4 | 77.2 | -4.2 | 3.5 | |
| PASS | 81.4 | 80.9 | -0.5 | 3.4 |
4.4 Transferability of Learned Sparse Structure
Inspired by studies suggesting the transferability of subnetworks between tasks [95, 96]. We investigate the transferability of PASS by posing two questions:() Can the sparse channel masks, learned in one task, be effectively transferred to other tasks? () Is the hypernetwork, once trained, applicable to other tasks? To answer Question (), we test the accuracy of subnetworks found on Tiny-ImageNet when fine-tuning on CIFAR- and CIFAR- and a pre-trained ResNet-18. To answer Question (), we measure the accuracy of the subnetwork finetuning on the target datasets, i.e., CIFAR- and CIFAR-. This subnetwork is obtained by applying hypernetworks, trained on Tiny-ImageNet, to the visual prompts of the respective target tasks. The results are reported in Table 1. We observe that the channel mask and the hypernetwork, both learned by PASS, exhibit significant transferability on target datasets, highlighting their benefits across various subsequent tasks. More interestingly, the hypernetwork outperforms transferring the channel mask in most target tasks, providing two hints: ❶ Our learned hypernetworks can sufficiently capture the important topologies in downstream networks. Note that there is no parameter tuning for the hypernetworks and only with an adapted visual prompt. ❷ The visual prompt can effectively summarize the topological information from downstream neural networks, enabling superior sparsification.

### v2

@@ Abstract
1 Introduction
Recently, large-scale neural networks, particularly in the field of vision and language modeling, have received upsurging interest due to the promising performance for both natural language [4, 10] and vision tasks [11, 2]. While these models have delivered remarkable performance, their colossal model size, coupled with their vast memory and computational requirements, pose significant obstacles to model deployment. To solve this daunting challenge, model compression techniques have re-gained numerous attention [12, 41] and knowledge distillation can be further adopted on top of them to recover optimal performance [19, 38]. Among them, model pruning is a well-established method known for its capacity to reduce model size without compromising performance and structural model pruning has garnered significant interest due to its ability to systematically eliminate superfluous structural components, such as entire neurons, channels, or filters, rather than individual weights, making it more hardware-friendly [23, 13].
In the context of structural pruning for vision models, the paramount task is the estimation of the importance of each structure component, such as channel or filters. It is a fundamental challenge since it requires dissecting the neural network behavior and a precise evaluation of the relevance of individual structural sub-modules. Previous methodologies [31, 13] have either employed heuristics or developed learning pipelines to derive scores, achieving notable performance. Recently, the prevailingness of natural language prompts has facilitated an emerging wisdom that the success of AI is deeply rooted in the quality and specificity of data that is originally created by human [46]. Techniques such as in-context learning [7] and prompting [36] have been developed to create meticulously designed prompts or input templates to escalate the output quality of LLMs. These strategies bolster the capabilities of LLMs and consistently achieve notable success across diverse downstream tasks. This offers a brand new angle for addressing the intricacies of structural pruning on importance estimation of vision models: How can we leverage the potentials within the input space to facilitate the dissection of the relevance of each individual structural component across layers, thereby enhancing structural sparsity?
One straightforward approach is directly editing input through visual prompt to enhance the performance of compressed vision models [42]. The performance upper bound of this approach largely hinges on the quality of the sparse model achieved by pruning, given that prompt learning is applied post-pruning. Moreover, when pruning is employed to address the intricate relevance between structural components across layers, the potential advantages of using visual prompts are not taken into consideration.
@@ 1 Introduction
Moreover, the complexities associated with inherent channel dependencies render the generation of sparse channel masks a challenging task. Due to this reason, many previous arts of pruning design delicate pruning metrics to recognize sparse subnetworks with smooth gradient flow [39]. To better handle the channel dependencies across layers during channel pruning, we propose to learn sparse masks using a recurrent mechanism. Specifically, the learned sparse mask for the recent layer largely depends on the mask from the previous layer in an efficient recurrent manner, and all the masks are learned by incorporating the extra information provided by visual prompts. The PASS framework is shown in Figure 1. Our contributions are summarized as follows:
@@ 2 Related Work
Structural Network Pruning. Structural pruning achieves network compression through entirely eliminating certain superfluous components from the dense network. In general, structural pruning follows three steps: (i) pre-training a large, dense model; (ii) pruning the unimportant channels based on criteria, and (iii) finetuning the pruned model to recover optimal performance. The primary contribution of various pruning approaches is located in the second step: proposing proper pruning metrics to identify the importance of channels. Some commonly-used pruning metric includes but not limited to weight norm [22], Taylor expansion [33], feature-maps reconstruction error [15] KL-divergence [32], greedy forward selection with largest loss reduction [43], feature-maps discriminant information [17].
@@
Prompting. In the realm of natural language processing, prompting has been acknowledged as an effective strategy to adapt pre-trained models to specific tasks [28]. The power of this technique was highlighted by GPT-3’s successful generalization in transfer learning tasks using carefully curated text prompts [4]. Researchers have focused on refining text prompting methods [37] and developed a technique known as Prompt Tuning. This approach involves using prompts as task-specific continuous vectors optimized during fine-tuning [25], offering comparable performance to full fine-tuning with a significant reduction in parameter storage and optimization. Prompt tuning’s application in the visual domain has seen significant advancement recently. Pioneered by [1], who introduced prompt parameters to input images, the concept was expanded by [6] to envelop input images with prompt parameters. [21] took this further, proposing visual prompt tuning for Vision Transformer models. Subsequently, [30] designed a prompt adapter to enhance these prompts. Concurrently, [45] integrated visual and text prompts in vision-language models, boosting downstream performance.
@@
Hypernetwork. Hypernetworks represent a specialized form of network architecture, specifically designed to generate the weights of another Deep Neural Network (DNN). This design provides a meta-learning approach that enables dynamic weight generation and adaptability, which is crucial in scenarios where flexibility and learning efficiency are paramount. Initial iterations of hypernetworks, as proposed by [47], were configured to generate the weights for an entire target DNN. While this approach is favorable for smaller and less complex networks, it constrains the efficacy of hypernetworks when applied to larger and more intricate ones. To address this limitation, subsequent advancements in hypernetworks have been introduced, such as the component-wise generation of weights [48] and chunk-wise generation of weights [5]. Diverging from the initial goal of hypernetworks, our work employs them to fuse visual prompts and model information for generating sparse channel masks.
@@ 3 PASS: Visual Prompt Locates Good Structure Sparsity
Notations. Let us consider a CNN with layers, and each layer contains its corresponding weight tensor , where {, , and } are the number of output/input channels and convolutional kernel size, respectively. The entire parameter space for the network is defined as . Similarly, a layer-wise binary mask is represented by , where “"/“" indicates removing/maintaining the associated channel. denotes our visual prompts. denotes the data of a target task.
Rationale. In the realm of structural pruning for deep neural networks, one of the key challenges is how to derive channel-wise importance scores for each layer. Conventional mechanisms estimate the channel significance either in a global or layer-wise manner [16, 13], neglecting the sequential dependency between adjacency layers. Meanwhile, the majority of prevalent pruning methods are designed in a model-centric fashion [13]. In contrast, an ideal solution to infer the high-quality sparse mask for one neural network layer should satisfy several conditions as follows:
@@ 3.1 Innovative Data-Model Co-designs through A Recurrent Hypernetwork
Modeling the Layer Sequential Dependency. The recurrent hyper-network in PASS adopts a Long Short-Term Memory (LSTM) backbone since it is particularly suitable for capturing sequential dependency. It enables an “auto-regressive" way to infer the structural sparse mask. To be specific, the LSTM mainly utilizes the previous layer’s mask , the current layer’s weights , and a visual prompt as follows:
where , the visual prompt provides an initial hidden state for the LSTM hyper-network, is the parameters of the LSTM model, and is the extra encoder to map the visual prompt into an embedding space. The channel-wise sparse masks () generated from the hyper-network are utilized to prune the weights of each layer as expressed by . represents the pruning of in-channels while denotes the pruning of out-channels.
@@
For achieving an optimal non-uniform layer-wise sparsity ratio, we adopt global pruning [18] that eliminates the channels associated with the lowest score values from all layers during each optimization step. This approach is grounded in the findings of [18, 13], which demonstrate that layer-wise sparsity derived using this method surpasses other extensively researched sparsity ratios.
@@ 4.1 Implementation Setups
Architectures and Datasets. We evaluate PASS using four traditional pre-trained models: ResNet-, ResNet-, ResNet-, and VGG- without BatchNormD and three advanced models:ResNeXt-50,ViT-B/16 and Swin-T, all pre-trained on ImageNet-1K. Our evaluation contains six target tasks: Tiny-ImageNet, CIFAR-, DTD, StanfordCars, and Food.
Baselines. We select five popular structural pruning methods as our baselines: () Group-L structural pruning [23] reduces the network channels via regularization. () GrowReg [40] prunes the network channels via regularization with a growing penalty scheme. () Slim [31] imposes channel sparsity by applying regularization to the scaling factors in batch normalization layers. () DepGraph [13] models the inter-layer dependency and group-coupled parameters for pruning and () ABC Pruner [26] performs channel pruning through automatic structure search.
Training and Evaluation. We utilize off-the-shelf models from Torchvision 111https://pytorch.org/vision/stable/index.html as the pre-trained models. During the pruning phase, we employ the SGD optimizer for the visual prompt, while the AdamW optimizer is used for the visual prompt encoder and the LSTM model for generating channel masks. Regarding the baselines, namely Group-L structural pruning, GrowReg, Slim, and DepGraph, they are trained based on this implementation 222https://github.com/VainF/Torch-Pruning and ABC Prunner is trained based on their official public code 333https://github.com/lmbxmu/ABCPruner. During the fine-tuning phase, all pruned models, inclusive of those from PASS and the aforementioned baselines, are fine-tuned with the same hyper-parameters. For all experiments, we report the accuracy of the downstream task during testing and the floating point operations (FLOPs) for measuring the efficiency.
@@ 4.2 PASS Finds Good Structural Sparsity
From Figure 2, several salient observations can be drawn: ❶ PASS consistently demonstrates superior accuracy across varying FLOPs values for all six evaluated downstream tasks. On one hand, PASS achieves higher accuracy under the same FLOPs. For example, it achieves higher accuracy than baselines under FLOPs among all the datasets. On the other hand, PASS attains higher speedup444Following [13], we report the theoretical speedup ratios and it is defined as in achieving comparable accuracy levels. For instance, to reach accuracy levels of , , and on CIFAR, StanfordCars, and Food respectively, the PASS method consistently realizes a speedup of at least ( VS ), outperforming the most competitive baseline. This consistent performance highlights the robustness and versatility of the PASS method across diverse scenarios. ❷ In terms of resilience to pruning, PASS exhibits a more gradual reduction in accuracy as FLOPs decrease. This trend is notably more favorable when compared with the sharper declines observed in other baseline methods. ❸ Remarkably, at the higher FLOPs levels, PASS not only attains peak accuracies but also surpasses the performance metrics of the fully fine-tuned dense models. For instance, PASS excels the fully fine-tuned dense models with {, , } on CIFRAR, DTD and FOOD datasets.
@@
Superior Performance across Model Architectures. We further evaluate the performance of PASS across multiple model architectures, namely VGG- without batch normalization 555The baseline Slim [31] is not applicable to this architecture., ResNet-, and ResNet- and compare it with the baselines. The results are shown in Figure 3. We observe that our PASS achieves a competitive performance across all architectures, often achieving accuracy close to or even surpassing the dense models while being more computationally efficient. For instance, To achieve an accuracy of on Tiny-ImageNet using ResNet-/ResNet- and accuracy using VGG-, our PASS requires fewer FLOPs compared to the most efficient baseline. These observations suggest that PASS can effectively generalize across different architectures, maintaining a balance between computational efficiency and model performance.
4.3 Experiments on ImageNet and Advanced Architectures
To draw a solid conclusion, we further conduct extensive experiments on a large dataset ImageNet using advanced pre-trained models such as ResNeXt-50, Swin-T, and ViT-B/16. The results are shown in Table 1. We observe that our method PASS demonstrates a significant speed-up with minimal accuracy loss, as indicated by the Acc., which is superior to existing methods like SSS [20], GFP [27], and DepGraph [13]. the resulting empirical evidence robustly affirms the effectiveness of PASS across both advanced neural network architectures and large-scale datasets.
4.4 Transferability of Learned Sparse Structure
Inspired by studies suggesting the transferability of subnetworks between tasks [8]. We investigate the transferability of PASS by posing two questions:() Can the sparse channel masks, learned in one task, be effectively transferred to other tasks? () Is the hypernetwork, once trained, applicable to other tasks? To answer Question (), we test the accuracy of subnetworks found on Tiny-ImageNet when fine-tuning on CIFAR- and a pre-trained ResNet-18. To answer Question (), we measure the accuracy of the subnetwork finetuning on the target datasets, i.e., CIFAR-. This subnetwork is obtained by applying hypernetworks, trained on Tiny-ImageNet, to the visual prompts of the respective target tasks. The results are reported in Table 2. We observe that the channel mask and the hypernetwork, both learned by PASS, exhibit significant transferability on target datasets, highlighting their benefits across various subsequent tasks. More interestingly, the hypernetwork outperforms transferring the channel mask in most target tasks, providing two hints: ❶ Our learned hypernetworks can sufficiently capture the important topologies in downstream networks. Note that there is no parameter tuning for the hypernetworks and only with an adapted visual prompt. ❷ The visual prompt can effectively summarize the topological information from downstream neural networks, enabling superior sparsification.
Arch. Method Base Pruned Acc. FLOPs ResNeXt-50 ResNeXt-50 77.62 - - 4.27 SSS [20] 77.57 74.98 -2.59 2.43 GFP [27] 77.97 77.53 -0.44 2.11 DepGraph [13] 77.62 76.48 -1.14 2.09 Ours (PASS) 77.62 77.21 -0.41 2.01 ViT-B/16 ViT-B/16 81.07 - - 17.6 DepGraph [13] 81.07 79.17 -1.90 10.4 Ours(PASS) 81.07 79.77 -1.30 10.7 Swin-T Swin-T 81.4 - - 4.49 X-Pruner [44] 81.4 80.7 -0.7 3.2 STEP [24] 81.4 77.2 -4.2 3.5 Ours(PASS) 81.4 80.9 -0.5 3.4